The units of an item available for sale during the year were as follows:

Jan. 1   Inventory 1,100 units @ $122
Feb. 17   Purchase 1,385 units @ $123
Jul. 21   Purchase 1,700 units @ $126
Nov. 23   Purchase 1,145 units @ $127

There are 1,220 units of the item in the physical inventory at December 31. The periodic inventory system is used. Do not round intermediate calculation and round final answer to nearest whole value.

a. Determine the inventory cost by the first-in, first-out method.
$

b. Determine the inventory cost by the last-in, first-out method.
$

c. Determine the inventory cost by the weighted average cost method.
$
Answer #1
Unit Unit Cost Total Cost
Beginning inventory 1100 122 134200
Purchase 1 1385 123 170355
Purchase 2 1700 126 214200
Purchase 3 1145 127 145415
Total 5330 664170

a) FIFO = (145415+75*126) = 154865

b) LIFO = (134200+120*123) = 148960

c) Weighted average = 664170/5330*1220 = 152024